OpenStudy (anonymous):

The revenue function for a school group selling n bookmarks is given by R(n) = 2n, and the total cost function is given by C(n) = 144 + 0.80n. Determine the number of bookmarks that need to be sold for the school group to break even.

OpenStudy (anonymous):

well take R(n)=C(n) and calculate n

OpenStudy (anonymous):

2n=144+.8n 1.2n=144 n=144/1.2 n=120

OpenStudy (anonymous):

for breakeven u need to sell 120 bookmarks :)

OpenStudy (anonymous):

Thank you!!!!!!! I was confused on where to start.

OpenStudy (anonymous):

u r welcome .. any time